OpenStudy (anonymous):

the population increase of iraq is given by the equation A=Aoe^0.028t,where t is the years after 2003 was 25 million round to the nearest million

OpenStudy (anonymous):

A=Aoe^0.028t, if A(0)=25M at 2003, t=2011-2003=8 years A=(25M)e^0.028*8 =(25M)e^.224 =31,276,775.49 by using calc or =31.277Million approximation